OpenStudy (anonymous):

Finding antiderivative of 8t^2+t/10 I know 8/3t^3 + ? + C not sure how to do the t/10 part.

OpenStudy (anonymous):

See in case of just take the constant term out of it and just integral the variable term.. As here 10 is constant so pull it out, and just take integral of t..

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

\[\int\limits \frac{t}{10}.dt \implies \frac{1}{10} \int\limits t.dt\]

OpenStudy (anonymous):

1/5 t^2

OpenStudy (anonymous):

Thanks!